OpenStudy (anonymous):

f(x)=2x^2-20x+4 Find the vertex

OpenStudy (anonymous):

i think i have the answer 5,-49 but i want someone to check

OpenStudy (anonymous):

(5, -46)

OpenStudy (anonymous):

how did you get 46?

OpenStudy (anonymous):

if you substitute 5 in for x -> 2(5)^2-20(5)+4 = 2(25)-100+4 = 50-100+4 = -50+4 = -46